A small, 2.00-mm-diameter circular loop with R = 1.00×10−2 Ω is at the center of a large 100-mm-diameter circular loop. Both loops lie in the same plane. The current in the outer loop changes from +1.0A to −1.0A in 0.100 s .

What is the induced current in the inner loop?
Answer #1

Solution)

Here, radius of smaller circular loop, r=2 mm/2= 1*10^-3 m

Radius of larger loop, r1=100 mm/2=50*10^-3 m

We know, Induced Emf, E=-dphi/dt

Induced Current, I= E/R= -d(pi*r^2)*B/rdt

Magnetic Field at center of loop, B= uoI/2r1

Now, Induced Current, I=(pi*r^2)*(uo/2r1*R)*(di/dt)

I=pi*(1*10^-3)^2*((4pi*10^-7)/(2*50*10^-3)*(1*10^-2))*(-1-1/0.1)

Induced Current, I=7.88*10^-8 A

=======